OpenStudy (anonymous):

Solve for x: 1) 4^2x-1=8^x 2) 27^2x-1=9^-x 3) (x+5)(3x-2)(23x)=0

OpenStudy (anonymous):

for 1 and 2: whenever you have x^a = x^b, you know that a = b In Q1: try converting 4 and 8 into 2^2 and 2^3 and see if you can do it.

OpenStudy (campbell_st):

rewrite question 1 with the same base.... 2 \[(2^2)^{2x -1} = (2^3)^x\] apply the appropriate index law than then equate the powers.